what is the origin of electrostatic forces.The origin of gravitational forces was explained by einstein by space time curvature.so does electrostatic force also originate from such type of curvature.if so then what type of curvature

Grade:11

1 Answers

Aravind Bommera
11 years ago

The fundamental equation of electrostatics is Coulomb''s law, which describes the force between two point charges. The magnitude of the electrostatic force between two point electric charges Q_1 and Q_2 is directly proportional to the product of the magnitudes of each charge and inversely proportional to the surface area of a sphere whose radius is equal to the distance between the charges:

F = \frac{Q_1Q_2}{4\pi r^2\varepsilon_0}\ ,

where ε0 is a constant called the vacuum permittivity or permittivity of free space, a defined value:

 \varepsilon_0 \ \stackrel{\mathrm{def}}{=}\ \frac {1}{\mu_0 {c_0}^2} = 8.854\ 187\ 817\ \times 10^{-12}   in A2s4 kg-1m−3 or C2N−1m−2 or F m−1.




Einstein proposed that spacetime is curved by matter, and that free-falling objects are moving along locally straight paths in curved spacetime. These straight paths are called geodesics. Like Newton''s first law of motion, Einstein''s theory states that if a force is applied on an object, it would deviate from a geodesic. For instance, we are no longer following geodesics while standing because the mechanical resistance of the Earth exerts an upward force on us, and we are non-inertial on the ground as a result. This explains why moving along the geodesics in spacetime is considered inertial.

Einstein discovered the field equations of general relativity, which relate the presence of matter and the curvature of spacetime and are named after him. The Einstein field equations are a set of 10 simultaneous, non-linear, differential equations. The solutions of the field equations are the components of the metric tensor of spacetime. A metric tensor describes a geometry of spacetime. The geodesic paths for a spacetime are calculated from the metric tensor.



May be it is false,electrostatic force is not originate from curvature.